# Multi-qubit Dynamical Decoupling for Enhanced Crosstalk Suppression

### v2

@@ Appendix C Noisy Simulation Results
We utilize a noise simulator to perform a T2 Ramsey experiment, aiming at demonstrating the effectiveness of staggered DD in mitigating the effects of $ZZ$ crosstalk. The noise model exclusively contains thermal relaxation without any gate or readout errors. We set the induced oscillation frequency to $100$ kHz to enable more precise measurement.
The procedure begins by inserting X2pm along with its staggered counterpart into the delay of the Ramsey experiment. Within these DD-inserted circuits, we incorporate an additional $ZZ(\theta)$ gate prior to each delay, with $\theta=\omega t$, serving to simulate $ZZ$ crosstalk across time evolution $t$. We evaluate the discrepancies in detuning frequency when inserting either standard or staggered DD sequences into Ramsey experiment circuits. This comparison is performed under two conditions: with and without the introduction of $ZZ$ crosstalk. The resulting detuning frequencies are detailed in table 3. Staggered DD is shown to be highly effective in suppressing the impact of $ZZ$ crosstalk, evident in its ability to maintain a consistent detuning frequency in comparison to the non-crosstalk scenario. Conversely, the presence of $ZZ$ crosstalk instigates a frequency shift when inserting standard DD. This noise simulation outcome aligns with the theoretical analysis, validating the capability of staggered DD in canceling out $ZZ$ crosstalk.
Table 3: T2 Ramsey experiment.
Experimental setup Detuning frequency (KHz)

### v3

@@ Appendix C Noisy Simulation Results
We utilize a noise simulator to perform a T2 Ramsey experiment, aiming at demonstrating the effectiveness of staggered DD in mitigating the effects of $ZZ$ crosstalk. The noise model exclusively contains thermal relaxation without any gate or readout errors. The thermal relaxation model includes both relaxation and dephasing errors. We set the induced oscillation frequency to $100$ kHz to enable more precise measurement.
The procedure begins by inserting X2pm along with its staggered counterpart into the delay of the Ramsey experiment. Within these DD-inserted circuits, we incorporate an additional $ZZ(\theta)$ gate prior to each delay, with $\theta=\omega t$, serving to simulate $ZZ$ crosstalk across time evolution $t$. We evaluate the discrepancies in detuning frequency when inserting either standard or staggered DD sequences into Ramsey experiment circuits. This comparison is performed under two conditions: with and without the introduction of $ZZ$ crosstalk. The resulting oscillation frequencies are detailed in table 4. In the case of no static $ZZ$, the oscillation frequency is the same as the set detuning (of 100 kHz). Static $ZZ$ is modeled by introducing $R_{ZZ}$ rotations in the idle times of the circuits, where the $Z$ rotations will be in the positive direction since the spectator qubit is in the ground state. The effect of this rotation is to change the observed oscillation frequency of the Ramsey experiment in the case of X2pm DD, whereas the staggered X2pm returns the observed oscillation frequency to nearly the set value of 100 kHz. This noise simulation outcome aligns with the theoretical analysis, validating the capability of staggered DD in canceling out $ZZ$ crosstalk.
Table 4: Observed oscillation frequencies with and without $ZZ$ crosstalk for a T2 Ramsey experiment with a set detuning of 100 KHz.
Simulated model Oscillation frequency (kHz)